Group Selection Page HELP Controls Group ITDM Vacuum WP78 SPB SQS XO ToBeMigrated2XWiki ToBeMigrated2GitlabPages ToBeMigratedByGroup Migrated2Zulip Archived2PDF not2MigrateOrArchive
SCS SXP Migrated2Zulip Photon Commissioning Team XRO DOC HED Migrated2Zulip WP74 MID PSPO_XO XO_Operation SQS Migrated2Zulip FXE Migrated2Zulip DET Sample Environment SPB Migrated2Zulip
OPERATION-2025 OPERATION-2024 OPERATION-2023 OPERATION-2022 OPERATION-2021 OPERATION-2020 OPERATION-2019 SASE1-2018 SASE1-2017 Experiment data
  OPERATION-2023 logbook, Contact: F.Wolff-Fabris, A.Galler, L.Samoylova, H.Sinn  Not logged in ELOG logo
Message ID: 434     Entry time: 04 Nov 2023, 02:29
Author: Johannes Moeller 
Group: MID 
Subject: Shift summary 
  • Stable delivery of SASE at 18 keV. The intensity dropped slightly from 1mJ to 0.88mJ during the evening, but everything good.
  • fixed the laser pulse-on-demand + FSSS scanning macro
  • We take some images series of the laser drilling process for different laser fluences
  • We started to take first time resolved data
ELOG V3.1.4-7c3fd00